OpenStudy (haleyelizabeth2017):

Use a graphing calculator to solve the equation 3 tan 1/3Θ=8 in the interval from 0 to 2pi.

OpenStudy (anonymous):

first divide both sides by 3

OpenStudy (haleyelizabeth2017):

\[\tan \frac{ 1 }{ 3 } \theta =\frac{ 8 }{ 3 }\]eh?

OpenStudy (anonymous):

yep. now take tan^-1 of both sides

OpenStudy (haleyelizabeth2017):

\[\frac{ 1 }{ 3 }\theta=1.21\]

OpenStudy (anonymous):

yep, now multiply both sides by 3

OpenStudy (haleyelizabeth2017):

\(\theta=3.64\)

OpenStudy (haleyelizabeth2017):

Wait...is that it? O.o

OpenStudy (anonymous):

yep 3.64 rad or 208.3 degrees

OpenStudy (haleyelizabeth2017):

ah, okay. Thank you :)